# Predicting New Research Directions in Materials Science using Large Language Models and Concept Graphs

### v1

@@ Abstract
Due to an exponential increase in published research articles, it is impossible for individual scientists to read all publications, even within their own research field. In this work, we investigate the use of large language models (LLMs) for the purpose of extracting the main concepts and semantic information from scientific abstracts in the domain of materials science to find links that were not noticed by humans and thus to suggest inspiring near/mid-term future research directions. We show that LLMs can extract concepts more efficiently than automated keyword extraction methods to build a concept graph as an abstraction of the scientific literature. A machine learning model is trained to predict emerging combinations of concepts, i.e. new research ideas, based on historical data. We demonstrate that integrating semantic concept information leads to an increased prediction performance. The applicability of our model is demonstrated in qualitative interviews with domain experts based on individualized model suggestions. We show that the model can inspire materials scientists in their creative thinking process by predicting innovative combinations of topics that have not yet been investigated.
@@
Promising new research directions often arise from combining research topics that have previously not been investigated together [uzzi_atypical_2013]. While experienced scientists possess vast domain knowledge enabling them to thoroughly explore research topics within (and adjacent to) their area(s) of expertise, finding new connections between their research topics and other yet unfamiliar topics to foster new ideas and findings is inherently challenging. Machine learning (ML) methods can help to look beyond the personal area of expertise by identifying previously unthought-of combinations of research topics, and thus enable the exploration of a vast hypothesis space beyond human intuition [varshney_big_2013, pinel_culinary_2015].
@@
Scientific information is contained in a plethora of research publications in a rich but unstructured manner, and this lack of structured information poses challenges for automated analysis [dunn_structured_2022, evans_advancing_2011]. Focusing on the extensive domain of material science, we first address the question of systematically extracting the main concepts of scientific articles, i.e. keywords or key phrases. Recent breakthroughs in Natural Language Processing (NLP) now allow the extraction of structured data from text and thereby enable the subsequent automated processing of this rich data [vaswani_attention_2017, brown_language_2020, torfi_natural_2021, kojima_large_2023, bubeck_sparks_2023, norouzi_conexion_2025]. Here, we investigate whether large language models (LLMs) can offer improvements over traditional algorithmic methods in this extraction process.
@@
After the identification and extraction of the concepts and their connections (i.e. the co-occurrence in the same article), we secondly investigate how to use this information to predict new combinations of topics. In a prior study, Krenn et al. proposed SemNet, a graph that tracks the evolution of scientific literature in the domain of Quantum Physics [krenn_predicting_2020]. The nodes of the SemNet graph are concepts, i.e. keywords extracted from text, using an algorithm called RAKE in conjunction with some predefined rules [rose_automatic_2010]. Apart from analyzing emerging trends within SemNet, the authors exploit the evolution of the graph to predict its further development. To this end, they derive topological properties such as node degree and use them as input for a neural network (NN) to predict future connections. In a Kaggle challenge, the participants predicted changes in a SemNet built from AI literature [krenn_predicting_2022]. While the most successful models combined specific hand-selected network features with machine learning techniques, such as NNs or Graph Neural Networks (GNNs), other participants employed purely theoretical or end-to-end ML approaches. All models of the participants could, however, only use the structure of SemNet as the real meaning behind the nodes was not revealed in the challenge.
@@
In this study, the information on materials science concepts is similarly compressed into what we call a concept graph. Given the advances in language encoder models [devlin_bert_2019, tenney_bert_2019], we further use the MatSciBERT model [gupta_matscibert_2021] to enrich the topological information of the nodes by additional information on the concepts in the form of semantic embeddings. After this, we further explore how the time evolution of this representation of the literature can be utilized by ML methods to perform link prediction.
In contrast to Schmidhuber who analyzes understanding, intelligence, and creativity in general and tries to evoke that in machines [schmidhuber_artificial_2010], we hope to foster creativity in humans through the assistance of AI by providing material scientists with a tool to propose new research directions, i.e. research topics holding great potential for synergies in combination with their own research area. To explore the real-world applicability of our model and its suggestions, we conduct interviews with researchers from the materials sciences. In these interviews, researchers assess how well suggested concepts generated by our model align with concepts from their own research.
Methods
The key steps of our approach are depicted in Figure 1. After gathering the abstracts of a large number of research publications in the domain of materials sciences, we extracted the main concepts, i.e. short key phrases consisting of few words from these abstracts, and used them as nodes in a concept graph that mirrors the (time-dependent) connectivity of the materials science concepts in literature. The final step of our workflow consists of performing link prediction on this graph based on both network properties, e.g. connectivity information, and semantic knowledge about the concepts captured in aggregated word embeddings.
Figure 1: Overview of the link prediction workflow: (a) gathering of materials science abstracts with OpenAlex [priem_openalex_2022], (b) extraction of concepts using LLMs, (c) creation of the semantics-aware-concept graph, and (d) prediction of new research directions.
Dataset
We prepared a dataset of published papers related to materials science. Data was obtained from OpenAlex by querying all publications listed at materials science-related journals, conferences, and other venues [priem_openalex_2022]. The retrieved papers were filtered based on language, length, and whether they have an abstract. For each publication, the title and abstract were cleaned and concatenated. Chemical formulae were extracted, stored separately, and later merged with the extracted concepts. The resulting dataset comprises approximately 221,000 articles published between 1955 and 2022, with relevant attributes being "title", "abstract", and "publication date". A more detailed description of the dataset generation is given in Section Dataset of the SI.
Concept extraction
Previous work used RAKE for concept extraction in conjunction with manual filtering to remove errors, i.e. phrases that do not represent semantic information, introduced by the imperfect statistical analysis performed in RAKE [krenn_predicting_2020, krenn_predicting_2022, rose_automatic_2010]. Instead, we opted for extracting concepts using fine-tuned LLMs (see Figure 2). To create a dataset for fine-tuning, 100 randomly chosen abstracts were first manually annotated by extracting and partially adjusting or even paraphrasing relevant and meaningful concepts as a preliminary step. We note that manual annotation is particularly sensitive to the labeler, as there is no unique way of extracting and defining concepts. Subsequently, we fine-tuned the LLaMa-2-13B base model [touvron_llama_2023, touvron_llama_2023-1] on our manually annotated abstracts for 4 epochs, using a learning rate of $5\cdot 10^{-4}$ and a batch size of 1 (see Section LLM Fine-tuning of the SI for details. Llama-2 models were SOTA when the work was performed in 2023, but will be replaced with newer models in future iterations of this work). The size of the model is a trade-off between accuracy and cost, as it is the largest model that can process 20 abstracts at once on an A100 GPU with 80GB of VRAM. To accelerate training and especially inference, we incorporated 8-bit quantization [dettmers_llmint8_2022] and low-rank adaptation (LoRa) techniques [hu_lora_2021, dettmers2023qlora] using Hugging Face’s PEFT module [peft].
Similar to Dunn et al.’s assisted annotation process [dunn_structured_2022], the fine-tuned model’s outputs were compared in the third step to the concepts extracted by GPT-3.5 [brown_language_2020] to efficiently identify and correct common mistakes of the fine-tuned model. In this way, we labeled 100 additional abstracts, and the base model was again fine-tuned on the larger dataset of 200 abstracts. The process of iteratively adding more automatically extracted and manually corrected concepts to fine-tune the model with more datapoints could, in principle, be repeated more often, but 200 labeled abstracts were enough for our use case. Finally, the resulting model was employed to extract concepts from all approximately 221,000 abstracts in our dataset, requiring approximately 160 GPU-hours. After extraction, we conducted minor post-processing of the extracted concepts by removing the remaining plural forms.
Concept graph
Figure 3: Example for a concept graph derived from two articles published 2012 (orange) in and 2015 (blue) with one overlapping concept. Possible new edges are marked as dotted gray lines together with their predicted probability of formation in the years 2017 to 2019.
To enrich the nodes of our concept graph with semantic information, we calculate concept embeddings and use them as node features. The procedure of calculating concept embeddings using MatSciBERT [gupta_matscibert_2021] is summarized in Figure 4: First, both the entire abstract and the previously extracted concepts are tokenized, and the embeddings are then calculated for the tokenized abstract. The next step consists of locating all instances of a concept in the abstract and averaging the embeddings of the tokens corresponding to the concept. For example, the concept ‘mechanical stress’ is tokenized as [4487, 1893], and the concept embedding is calculated as the average of the corresponding representations in the embedded abstract, i.e. those corresponding to the positions of the sequence [4487, 1893] in the tokenized abstract. To derive a singular representation for each concept per abstract, we average the embeddings of all its occurrences. In cases where a concept does not appear verbatim in an abstract — e.g. due to the normalization processes during the initial concept extraction — we take the mean embedding of all tokens (while excluding the start and end token in the abstract as its representation). As the final step, the average embedding for identical concepts across different abstracts is calculated to obtain a single embedding for each concept, and thus for each node.
Link prediction
The previous method used by Krenn et al. to predict new links in their concept graph exclusively relied on abstract local graph properties, either through a purely graph-theoretical approach using hand-crafted features in conjunction with machine learning or through employing end-to-end machine learning methods [krenn_predicting_2020, krenn_predicting_2022]. Here, we investigate whether integrating semantic knowledge about the concepts can improve link prediction. In particular, we use concept embeddings, i.e., high-dimensional vectors that capture semantic information, to make the semantic information integrable into the link prediction task.
Given the concept graph $G$, we treat link prediction as a binary classification task. The objective of the ML model is thus to predict whether a new edge is formed between an arbitrary pair of previously unconnected vertices $(u,v)$ in the time range $T=[T_{\mathrm{start}},T_{\mathrm{end}}]$. We chose $T_{\mathrm{start,train}}=2017$ and $T_{\mathrm{end,train}}=2019$ for training, i.e. our model has access to the entire data up to and including 2016 while its predictions are made for the years 2017, 2018, and 2019. We illustrate link prediction on a rudimentary concept graph in 3.
The prediction task has an inherent strong label imbalance since the likelihood of a randomly selected vertex pair $(u,v)$ forming a link throughout three years is extremely low. While there are for example 18.7 billion possible new edges that could form between 2017 and 2019, only 1.3 million new edges ($0.007\%$) can actually be observed during this period. To address this imbalance, we oversampled positive labels by using a fixed percentage of positive examples per batch in the training process.
A modified version of Krenn et al.’s densely connected neural network [krenn_predicting_2022], which purely relies on graph properties at different points in time, was used as the Baseline model. Specifically, the degree of a node $u$ ($\sum_{i=1}^{n}A_{u,i}$) and the sum of all 2-length paths from $u$ ($\sum_{i=1}^{n}A_{u,i}^{2}$) were calculated for different years in the range $t=[T_{\text{start,train}}-5,T_{\text{start,train}}-1]$, where $A_{t}$ denotes the binary adjacency matrix of $G_{t}$. These features were then concatenated for a given pair of nodes $(u,v)$ to result in a 20-dimensional baseline feature vector. In the second Concept embeddings model, the concatenated concept embeddings of $u$ and $v$ were used instead as the (1,536-dimensional) feature vector for the neural network classifier, to test their information content and relevance for the link prediction task.
The Baseline model was then combined with the Concept embeddings model in two hybrid models, for the first of which (Combination of features), a concatenation of the feature vectors of the two previous models was used as the input. The hyperparameters of the three neural networks were optimized using a comprehensive grid search varying the number of neurons per layer, the percentage of positive samples in each batch, the learning rate, and the dropout probability (see Table S1 in the SI for a list of optimized hyperparameters).
The second hybrid model (Combination of models) uses a weighted output of the optimized Baseline and Concept embeddings models where an optimal weighting of 3:2 was determined using hyperparameter optimization. We note that there are many other possible hybrid models beyond concatenating the two input vectors and calculating the weighted average of the output probabilities. The two parts of the input could be passed through a first set of layers separately before the two outputs are concatenated and passed through a second set of layers. The optimization of the architecture of the NN was however outside the scope of this study, as our goal mainly consisted in showing that the inclusion of the concept embeddings improves link prediction.
@@ Results
Using an LLM-based approach (see Methods), approximately 510,000 chemical formulae and 3,600,000 concepts were extracted from the 221,000 abstracts in our database, which corresponds to an average of 2.3 chemical formulae and 16.3 concepts per abstract. The extracted concepts were then condensed to approximately 52,000 unique formulae and 1,241,000 unique concepts by removing duplicates. In general, our method resulted in more precise concept extraction than rule-based approaches (see Tables S2 - S7). Due to the extraction capabilities of LLMs, the amount of manual annotation work to generate the initially required data is negligible, especially as our iterative approach (Figure 2) reduces the manual effort to a minimum. Remarkably, the fine-tuned LLMs were able to extract concepts that were not exactly present in the text. Table 1 shows selected examples to demonstrate the capabilities of fine-tuned LLMs for nominalisation, removal of fill words such as ‘of’, plural-to-singular conversion, and formatting corrections.
@@
To construct a concept graph, we only include concepts that appeared at least 3 times and consist of at least 2 words. Due to this filtering, the resulting graph comprises approximately 137,000 nodes and 13,000,000 edges in total, making the calculation of topological features that require the squaring of the adjacency matrix possible in the first place. We present an overview of the 25 most frequently encountered concepts and formulas in Table S8.
@@
An analysis of the node degree distribution in the resulting graph shows that the majority of the nodes have a degree between 30 and 1,000 (Figure S2). While a few concept nodes act as hubs – having a notably larger number of connections than others – most of the concepts in the graph are directly linked to only a few others, making the resulting graph sparse. The evolution of the concept graph over time shows that connectivity further increases, as more papers are being published using already existing concepts. We observe an increase in concept centralization, i.e. that fewer and fewer nodes account for a larger share of the total connections (Figure S3).
We visualize all concepts by projecting their high-dimensional concept embeddings to 2 dimensions using UMAP [mcinnes_umap_2018]. We display the result which we call the ‘Map of materials science’ in Figure 5 (interactive version that can be explored on inspire.aimat.science). We then run nearest neighbor queries [scikit-learn] on the concept embeddings to explore whether these 768-dimensional vectors capture semantic meaning. The example queries listed in Table S9 show the striking similarity between the queried concept and its nearest neighbors.
@@
Figure 5: Map of the materials science. The map displays concepts we extracted from our abstract corpus. We took all concept embeddings and projected them in a 2D plane using UMAP [mcinnes_umap_2018]. For each square of length 2, the concept with the highest degree is highlighted and annotated (‘Highest degree in tile’).
@@
To statistically assess the performance of our link prediction models, we evaluate their performance on a held-out test set for edge formation in the period between 2020 and 2022, consisting of 2,000,000 node pairs. We then complement this with a qualitative analysis of the real-world applicability of the models based on human expert knowledge.
@@
Figure 6 (a) shows the ROC curves for the prediction of link formation during the test period, as they illustrate the capacities of the model to distinguish between classes across all possible classification thresholds. ROC curves are particularly useful for imbalanced datasets because they evaluate performance independently of class distribution [fawcett_roc_2004]. Although the Baseline model performs slightly better (AUC 0.9109) than the Concept embeddings model (AUC 0.8855), it already shows that the semantic information contained in the concept embeddings can be utilized by our model architecture. Furthermore, the performance of the two combination models demonstrates that the link prediction task benefits from incorporating semantic knowledge on top of local graph features: While the Combination of features model already shows a slightly improved AUC of 0.9147, the Combination of models approach exhibits a significant performance leap in the AUC metric, which reaches a maximum of 0.9372 when scaling the Baseline and Concept embeddings model predictions by 0.6 and 0.4, respectively. We speculate that gradient descent optimization on a unified feature vector (concatenating the features of the Baseline and Concept Embeddings model) – as it is done in the Combination of features model – might not be as effective as optimizing them individually. The distinct nature of baseline features versus high-dimensional concept embeddings could lead to the gradient for each batch becoming a suboptimal compromise between the gradients suited for each feature type in isolation.
@@
Figure 6: Performance metrics (ROC curves and the respective AUC) for our link prediction models on the test set ($T_{\text{test}}=[2020,2022]$). Markers highlight the performances at a threshold of $0.5$. (a) The panel shows the ROC curves on all data points with a zoomed-in view of the low false positive rate region in the inset. (b) and (c) The panels display the respective performance metrics for $d_{\text{prev}}=2$ and $d_{\text{prev}}=3$.
@@
We investigated the predicted new links with regard to the previous node distance $d_{\text{prev}}$, i.e. the shortest path distance between two nodes in $T_{\text{test}}=[2020,2022]$ before they become directly connected. An analysis of the graph shows its dense interconnectedness as certain prevalent concepts in materials science, like ‘mechanical property’ and ‘x ray diffraction’ (Table S8) have edges with the majority of nodes, which leads to a large number of short distances between many concept pairs. Although the graph consists of 137,000 nodes, nearly all of the unlinked concept pairs in the test set were already connected through one ($d_{\text{prev}}=2$, 43.3%) or two ($d_{\text{prev}}=3$, 56.5%) concepts. The distribution of $d_{\text{prev}}$ is even more biased towards short paths for the positive samples, i.e. the connections that actually formed during the test period. 290 of 307 emerging edges (94.5%) were found to have $d_{\text{prev}}=2$, while only 17 (5.5 %) had a previous distance of 3 which clearly shows that the proximity of two nodes in the concept graph increases the probability for a new edge to form between them. Samples at $d_{\text{prev}}=4$ were found to be very scarce (0.2% of the total samples, all negatives), and therefore excluded from further analysis.
While the Baseline model tends to correctly predict emerging edges primarily at a distance of 2 (212/213 TP have $d_{\text{prev}}=2$) with a recall of 73.1%, it performs much worse for $d_{\text{prev}}=3$ (recall: 5.9%). By contrast, the Concept embeddings model achieves a recall of 35.3% for d=3 while only slightly compromising on the recall at $d_{\text{prev}}=2$ (70.0%). The results are summarized in Table 2. Note that optimizing the classification metrics by changing the classification threshold of 0.5 for a positive prediction is outside the scope of this work, which mainly aims at rating non-existing links with respect to their future emergence rather than accurately predicting whether a new link will form or not.
In addition, we also separately calculated the ROC curves and the corresponding AUCs for both $d_{\text{prev}}=2$ and $d_{\text{prev}}=3$, and the results (Figure 6 (b) and (c)) emphasize the Baseline model’s failure to correctly categorize most positives with $d_{\text{prev}}=3$. This not only highlights the inherent challenge of predicting positives at greater distances but also indicates that the integration of semantic information enhances the model’s ability to forecast connections between concept pairs that are further apart in the graph. However, these emerging connections with larger previous node distances are particularly interesting and hold great potential to broaden the scientific scope beyond the more obvious new research directions.
@@
Table 2: True positives, false negatives, and recall (true positive rate) for the link predictions of the Baseline and Concept embeddings model split by previous node distances (full data in Table S10)
Model $d_{\text{prev}}$ TP FN Recall
Baseline 2 212 78 73.1%
3 001 16 05.9%
Concept embeddings 2 203 87 70.0%
3 006 11 35.3%
@@
As the second part of the model performance analysis, we furthermore conducted interviews with ten material scientists (human experts) based on individualized reports containing individualized recommendations for combinations of concepts suggested by our best prediction model (Combinations of models). The suggestions were subsequently discussed in the interviews to assess and clarify the proposed concept combinations. As the limited number of interviewees only allows for a qualitative analysis, all findings are anecdotal. Nonetheless, the expert feedback sheds light on the usefulness of the suggestions provided by our model.

### v2

@@ Abstract
Due to an exponential increase in published research articles, it is impossible for individual scientists to read all publications, even within their own research field. In this work, we investigate the use of large language models (LLMs) for the purpose of extracting the main concepts and semantic information from scientific abstracts in the domain of materials science to find links that were not noticed by humans and thus to suggest inspiring near/mid-term future research directions. We show that LLMs can extract concepts more efficiently than automated keyword extraction methods to build a concept graph as an abstraction of the scientific literature. A machine learning model is trained to predict emerging combinations of concepts, i.e. new research ideas, based on historical data. We demonstrate that integrating semantic concept information leads to an increased prediction performance. The applicability of our model is demonstrated in qualitative interviews with domain experts based on individualized model suggestions. We show that the model can inspire materials scientists in their creative thinking process by predicting innovative combinations of conceptsthat have not yet been investigated.
@@
Promising new research directions often arise from combining concepts that have previously not been investigated together [1]. While experienced scientists possess vast domain knowledge enabling them to thoroughly explore research topics within (and adjacent to) their area(s) of expertise, finding new connections between their research topics and other yet unfamiliar topics to foster new ideas and findings is inherently challenging. Machine learning (ML) methods can help to look beyond the personal area of expertise by identifying previously unthought-of combinations of research topics, and thus enable the exploration of a vast hypothesis space beyond human intuition [2, 3].
@@
Scientific information is contained in a plethora of research publications in a rich but unstructured manner, and this lack of structured information poses challenges for automated analysis [4, 5]. Focusing on the extensive domain of material science, we first address the question of systematically extracting the main concepts of scientific articles, i.e. keywords or key phrases. Recent breakthroughs in Natural Language Processing (NLP) now allow the extraction of structured data from text and thereby enable the subsequent automated processing of this rich data [6, 7, 8, 9, 10, 11]. Here, we investigate whether large language models (LLMs) can offer improvements over traditional algorithmic methods in this extraction process.
@@
After the identification and extraction of the concepts and their connections (i.e. the co-occurrence in the same article), we secondly investigate how to use this information to predict new combinations of concepts. In a prior study, Krenn et al. proposed SemNet, a graph that tracks the evolution of scientific literature in the domain of Quantum Physics [12]. The nodes of the SemNet graph are concepts, i.e. keywords extracted from text, using an algorithm called RAKE in conjunction with some predefined rules [13]. Apart from analyzing emerging trends within SemNet, the authors exploit the evolution of the graph to predict its further development. To this end, they derive topological properties such as node degree and use them as input for a neural network (NN) to predict future connections. In a Kaggle challenge, the participants predicted changes in a SemNet built from AI literature [14]. While the most successful models combined specific hand-selected network features with machine learning techniques, such as NNs or Graph Neural Networks (GNNs), other participants employed purely theoretical or end-to-end ML approaches. All models of the participants could, however, only use the structure of SemNet as the real meaning behind the nodes was not revealed in the challenge.
@@
In this study, the information on materials science concepts is similarly compressed into what we call a concept graph. Given the advances in language encoder models [15, 16], we further use the MatSciBERT model [17] to enrich the topological information of the nodes by additional information on the concepts in the form of semantic embeddings. After this, we further explore how the time evolution of this representation of the literature can be utilized by ML methods to perform link prediction.
Recent advances show that graph-based approaches can accelerate discovery in materials science: SciAgents employs multi-agent graph reasoning, Graph-PRefLexOR integrates symbolic graph abstractions with LLMs, and generative knowledge extraction with graph representations further supports hypothesis generation [18, 19]. Complementary efforts on AI-driven ideation include SciMuse and SciMON, which use enriched co-occurrence and temporal knowledge graphs for novel idea generation, ResearchAgent, which iteratively refines literature-grounded ideas with knowledge-augmented LLMs, and SCI-IDEA, which applies context-aware embeddings for systematic ideation [20, 21, 22, 23]. In contrast to approaches which analyze understanding, intelligence, and creativity in general and try evoke these in machines [24], we hope to foster creativity in humans through the assistance of AI by providing material scientists with a tool to propose new research directions, i.e. previously uncombined concept combinations. To explore the real-world applicability of our model and its suggestions, we conduct interviews with researchers from the materials sciences. In these interviews, researchers assess how well suggested concepts generated by our model align with concepts from their own research.
@@ Results
Using an LLM-based approach (see Methods), approximately 510,000 chemical formulae and 3,600,000 concepts were extracted from the 221,000 abstracts in our database, which corresponds to an average of 2.3 chemical formulae and 16.3 concepts per abstract. The extracted concepts were then condensed to approximately 52,000 unique formulae and 1,241,000 unique concepts by removing duplicates. In general, our method resulted in more precise concept extraction than rule-based approaches S1. Due to the extraction capabilities of LLMs, the amount of manual annotation work to generate the initially required data is negligible, especially as our iterative approach (see the Methods section below) reduces the manual effort to a minimum. Remarkably, the fine-tuned LLMs were able to extract concepts that were not exactly present in the text. Table 1 shows selected examples to demonstrate the capabilities of fine-tuned LLMs for nominalisation, removal of fill words such as ‘of’, plural-to-singular conversion, and formatting corrections.
@@
To construct a concept graph, we only include concepts that appeared at least 3 times and consist of at least 2 words. Due to this filtering, the resulting graph comprises approximately 137,000 nodes and 13,000,000 edges in total, making the calculation of topological features that require the squaring of the adjacency matrix possible in the first place. We present an overview of the 25 most frequently encountered concepts and formulas in Supplementary Table 1.
@@
An analysis of the node degree distribution in the resulting graph shows that the majority of the nodes have a degree between 30 and 1,000 (Supplementary Fig. 8). While a few concept nodes act as hubs – having a notably larger number of connections than others – most of the concepts in the graph are directly linked to only a few others, making the resulting graph sparse. The evolution of the concept graph over time shows that connectivity further increases, as more papers are being published using already existing concepts. We observe an increase in concept centralization, i.e. that fewer and fewer nodes account for a larger share of the total connections (Supplementary Fig. 9).
@@
Figure 1: Map of the materials science. UMAP[25] 2D projection of all extracted concepts with the highest-degree concepts in each square of length 2 highlighted and annotated (‘Highest degree in tile’). Yellow and purple background colors respectively indicate high and low concept densities calculated using Kernel Density Estimation (KDE) is indicated by background color.
We visualize all concepts by projecting their high-dimensional concept embeddings to 2 dimensions using UMAP [25] with default settings. We display the result which we call the ‘Map of materials science’ in Fig. 1 (an interactive version that can be explored on inspire.aimat.science). We then run nearest neighbor queries [26] on the concept embeddings to explore whether these 768-dimensional vectors capture semantic meaning. The example queries listed in Supplementary Table 2 show the striking similarity between the queried concept and its nearest neighbors.
@@
To statistically assess the performance of our link prediction models, we evaluate their performance on a held-out test set for edge formation in the period between 2020 and 2022, consisting of 2,000,000 node pairs and containing 307 (0.015%) positives, i.e. emerging edges. We then complement this with a qualitative analysis of the real-world applicability of the models based on human expert knowledge.
@@
Fig. 2 (a) shows the ROC curves for the prediction of link formation during the test period, as they illustrate the capacities of the model to distinguish between classes across all possible classification thresholds. ROC curves are particularly useful for imbalanced datasets because they evaluate performance independently of class distribution [27]. More information about test set creation, as well as detailed results (Precision/Recall@k) can be found in section S3. Although the Baseline model (a modified version of Krenn et al. [14]; see Methods) performs slightly better (AUC 0.9109) than the Concept embeddings (MatSciBERT) model (AUC 0.8855), the performance of the latter already shows that the semantic information contained in the concept embeddings can be used by our model architecture. A GNN model based on the GraphSAGE architecture (see Methods) surpasses the Baseline with an AUC of 0.9288, suggesting that while both models access the same input features, the GNN effectively leverages additional structural signals to improve performance. The Pure Text Baseline (implemented via a fine-tuned MatSciBERT [17]) also exploits this semantic information and performs similarly overall, though worse for $d_{\text{prev}}=3$ at a $5\times$ higher inference cost. Furthermore, the performance of the three hybrid models demonstrates that the link prediction task benefits from incorporating semantic knowledge on top of local graph features: While the Combination of features model already shows a slightly improved AUC of 0.9147, the Mixture of Baseline and Embeddings and Mixture of GNN and Embeddings models approach exhibits a significant performance leap in the AUC metric, which reaches a maximum of 0.9433 when scaling the GNN and Concept embeddings model predictions by 0.5 and 0.5, respectively. We speculate that gradient descent optimization on a unified feature vector (concatenating the features of the Baseline and Concept Embeddings model) – as it is done in the Combination of features model – might not be as effective as optimizing them individually. The distinct nature of baseline features versus high-dimensional concept embeddings could lead to the gradient for each batch becoming a suboptimal compromise between the gradients suited for each feature type in isolation. While MatSciBERT may under-represent emerging or interdisciplinary concepts, it still benefits from the base BERT knowledge, and tokenization ensures meaningful embeddings. In our experiments, MatSciBERT (AUC 0.8855) outperformed BERT (AUC 0.8547), indicating an advantage of domain-specific embeddings, though BERT still offers a reasonable baseline.
@@
Figure 2: Performance metrics (ROC and the respective AUC) for our link prediction models on the test set ($T_{\text{test}}=[2020,2022]$). Markers highlight the performances at a threshold of $0.5$. (a) ROC curves on all data points with a zoomed-in view of the low false positive rate region in the inset. Panels (b) and (c) display the respective performance metrics for $d_{\text{prev}}=2$ and $d_{\text{prev}}=3$. Best result in bold.
@@
We investigated the predicted new links with regard to the previous node distance $d_{\text{prev}}$, i.e. the shortest path distance between two nodes in $T_{\text{test}}=[2020,2022]$ before they become directly connected. An analysis of the graph shows its dense interconnectedness as certain prevalent concepts in materials science, like ‘mechanical property’ and ‘x ray diffraction’ (Supplementary Table 1) have edges with the majority of nodes, which leads to a large number of short distances between many concept pairs. Although the graph consists of 137,000 nodes, nearly all of the unlinked concept pairs in the test set were already connected through one ($d_{\text{prev}}=2$, 43.3%) or two ($d_{\text{prev}}=3$, 56.5%) concepts. The distribution of $d_{\text{prev}}$ is even more biased towards short paths for the positive samples, i.e. the connections that actually formed during the test period. 290 of 307 emerging edges (94.5%) in the test set were found to have $d_{\text{prev}}=2$, while only 17 (5.5 %) had a previous distance of 3 which clearly shows that the proximity of two nodes in the concept graph increases the probability for a new edge to form between them. Samples at $d_{\text{prev}}=4$ were found to be very scarce (0.2% of the total samples, all negatives), and therefore excluded from further analysis.
While the Baseline model tends to correctly predict emerging edges primarily at a distance of 2 (212/213 TP have $d_{\text{prev}}=2$) with a recall of 73.1%, it performs much worse for $d_{\text{prev}}=3$ (recall: 5.9%). By contrast, the Concept embeddings model achieves a significantly better recall of 35.3% (p < 0.05, DeLong test [28] for $d_{\text{prev}}=3$ while only slightly compromising on the recall at $d_{\text{prev}}=2$ (70.0%). Interestingly, the GNN model matches this performance at $d_{\text{prev}}=3$, demonstrating that improved structural processing can rival the benefits of semantic embeddings for distant connections. The results are summarized in 2. The high number of false positives, especially at $d_{\text{prev}}=3$ is not a problem in itself, as those combinations may remain scientifically plausible and will subsequently be evaluated by human scientists. Hence, we prioritize recall over precision in order not to miss valuable ideas.
@@
Table 2: Confusion matrix classes for the link predictions of the baseline, Concept Embeddings, GNN, and GNN-Mixture models split by previous node distances.
Positives Negatives
Model $d_{\text{prev}}$ TP FN Recall FP TN
Baseline 2 212 78 73.1% 115,807 0,751,021
3 001 16 05.9% 005,342 1,124,056
4 000 00 000,000 0,003,467
Concept embeddings 2 203 87 70.0% 120,575 0,746,253
3 006 11 035.3% 026,209 1,103,189
4 000 00 000,027 0,003,440
GNN 2 253 37 87.2% 223,637 0,643,191
3 003 14 017.6% 031,512 1,097,886
4 000 00 000,009 0,003,458
Mixture of GNN and Embeddings 2 248 42 85.5% 153,455 0,713,373
3 005 12 029.4% 011,753 1,117,645
4 000 00 000,006 0,003,461
Note that optimizing the classification metrics by changing the classification threshold of 0.5 for a positive prediction is outside the scope of this work, which mainly aims at rating non-existing links with respect to their future emergence rather than accurately predicting whether a new link will form or not.
In addition, we also separately calculated the ROC curves and the corresponding AUCs for both $d_{\text{prev}}=2$ and $d_{\text{prev}}=3$, and the results (Fig. 2 (b) and (c)) emphasize the Baseline model’s failure to correctly categorize most positives with $d_{\text{prev}}=3$. This not only highlights the inherent challenge of predicting positives at greater distances but also indicates that the integration of semantic information enhances the model’s ability to forecast connections between concept pairs that are further apart in the graph. However, these emerging connections with larger previous node distances are particularly interesting and hold great potential to broaden the scientific scope beyond the more obvious new research directions. Ultimately, the Mixture of GNN and Embeddings yields the highest AUC for these distant connections, outperforming the individual models by effectively combining structural and semantic signals.
We evaluated our Baseline model on the Science4Cast benchmark, where it achieved an AUROC of 0.9088, ranking second among all reported approaches in [14]. This demonstrates that a deep neural network trained on a large set of semantically meaningful features can outperform most competing methods, including those based on Common Neighbors and node2vec embeddings combined with a Transformer architecture [6, 29, 30]. As Science4Cast does not contain any semantic or text information about the meaning of nodes, we cannot apply and benchmark our embedding-based models on the Science4Cast challenge.
@@
As the second part of the model performance analysis, we furthermore conducted interviews with ten material scientists (human experts) based on individualized reports containing individualized recommendations for combinations of concepts suggested by our prediction model (Mixture of Baseline and Embeddings). Personalized recommendations were generated using the Mixture of Baseline and Embeddings model which is marginally less performant (<0.01<0.01 AUC) than the Mixture of GNN and Embeddings model, because the GNNs were only tested at a later stage of our study. The suggestions were subsequently discussed in the interviews to assess and clarify the proposed concept combinations. The small sample size, i.e. the small number of interviewees, as well as a potential selection bias limit the robustness of our conclusions and only allow a qualitative analysis and anecdotal findings. Nonetheless, the expert feedback sheds light on the usefulness of the suggestions provided by our model.